OpenStudy (anonymous):

evaluate the integral: ∫cos²x on (0, π/4)

OpenStudy (anonymous):

Integral of cos^2 is 1/2 ( x +sinxcosx ) So plug in for your limits of integration and solve!

OpenStudy (anonymous):

you need power reduction formula

OpenStudy (anonymous):

\[\cos^2x=\frac{1+\cos(2x)}{2}\]

OpenStudy (anonymous):

this is what DS did above i believe then he just took the double angle and use that